

********************************************************************
UARS WEEKLY STATUS REPORT - 07 JUL 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Our attempted observation of Mercury failed due to targeting
errors. We will try again in September.
Level 2 processing is halted due to bugs in the OA_SSPP_ATT and
OA_SAT_ATT services related to data gaps. CDHF and CSC personnel have
released engineering patches to these routines and we are beginning to
incorporate them into a test release of our production software.
***********
SUSIM
***********
Because of low power conditions near yaw around, SUSIM was switched off
on July 3. SUSIM's temperatures soon began to fall quickly, reaching 9
degrees on the 6th, before more power was applied to the heaters. The
temperature situation delayed the SUSIM power-on by a day. By July 7
when SUSIM powered on, the temperatures reached 11 degrees. Normal
operations will not begin until SUSIM has a temperature of 15 degrees.
Otherwise, SUSIM continues to gather a near normal set of daily solar
scans scheduled for periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM/AXIS-1 was activated on June 27 and deactivated on June 29 due to
spacecraft power depletion. The PEM/AXIS was not reactivated again before the
July YAW.
At the present time, PEM energy deposition is not being generated.
Recent data configuration changes have affected PEM software causing multiple
failures. These failures are being reviewed by the PEM team. In order to
correct for new telemetry requirements, PEM will need to update processing
code with a patch delivery. The delivery patch will not include changes which
would allow the PEM software to execute on a DEC Alpha machine.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the UARS Monthly Status Review
meeting held on 5 July 2000.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 24 June 2000 (Orbit 48018, GMT Day 176) through 30 June 2000
(Orbit 48121 , GMT Day 182).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 179/0018 48062 AXIS On
PEM 181/0241 48093 AXIS Off
HALOE 182/0522 47810 Off
The beta angle ranged between the angles of 31.0 degrees to 15.0 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
2.0 degrees on 04 July 2000 (DOY 186).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 23 June 2000 is provided in Appendix D.
The clock error ranged from -6.2 to 5.7 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 01
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 24
% HALOE = 04
% PEM = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/145 - 2287.497008
- Transponder B: 00/152 - 2287.498920
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/181 show the observatory data loss to be 296 hours, 32
minutes, 34.041 seconds (an increase of 12 minutes 26.328 seconds since the
last report period). This is a 0.40 percent data loss which equals a 99.60
percent data capture for the mission.
The increase of 12 minutes 26.328 seconds occurred over four (4) days:
- 0 minutes 57.264 seconds on Day 175 due data hits,
- 3 minutes 44.336 seconds on Day 177 due to a software problem at WSC,
- 7 minutes 28.416 seconds on Day 179 due to an unknown reason,
- 0 minutes 16.312 seconds on Day 180 due to data hits.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to test the new TMONs. Once the TMONs have been thoroughly
tested, they will be presented for review and approval before uplink.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
Per the request of SOLSTICE Operations, the SSPP was commanded (via stored
commands) to track the planet Mercury on GMT day 182, from 00:55:00 to 00:57:25
Z. Since the UARS Command Management System was designed to only formulate
solar and stellar SSPP tracking sessions, the SSPP had to be "tricked" into
tracking Mercury. Similar tracking sessions were performed in the past for
Jupiter, as well as the Hale Bopp comet.
The FOT continued to provide assistance to Brenda Coleman and Jim Kast of the
CMS test team. Brenda and Jim are still in the process of getting the Back-Up
CMS configured for testing. The FOT has been aiding the test team in
configuration issues, establishing a directory structure to mirror the
operational structure, and general procedural issues.
MEETINGS
The FOT attended the UARS FSW CCB on 26 June 2000. At this meeting it was
decided that 3 TDRS is the option with the highest possibility of success and
the least amount of risk to the mission. All efforts from FSW and the CMS
development team will be pursuant of this option.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for July 04, 2000 (DOY 186). This maneuver
will be a Reverse to Forward Flight maneuver (Southern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 30 June 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
48019 176/0333 00-292 NBTR PLAYBACK
48021 176/0650 00-292 NBTR NOOPS D-176
48027 176/1714 00-292 NBTR NOOPS D-176
48028 176/1756 00-293 NBTR NOOPS D-177
48033 177/0355 00-293 NBTR PLAYBACK
48036 177/0653 00-293 NBTR NOOPS D-177
48043 177/1758 00-294 NBTR NOOPS D-178
48049 178/0337 00-294 NBTR PLAYBACK
48051 178/0655 00-294 NBTR NOOPS D-178
48057 178/1620 00-295 HALOE HALOE FOV
48058 178/1800 00-296 HALOE HALOE FOV
48058 178/1800 00-297 HALOE HALOE FOV
48058 178/1800 00-297 NBTR NBTR NOOPS D-179
48062 179/0018 00-299 PEM AXIS1 Power ON
48064 179/0338 00-298 NBTR PLAYBACK
48065 179/0515 00-298 NBTR NOOPS D-179
48072 179/1720 00-298 NBTR NOOPS D-179
48073 179/1802 00-300 NBTR NOOPS D-180
48079 180/0339 00-300 NBTR PLAYBACK
48080 180/0515 00-300 NBTR NOOPS D-180
48088 180/1803 00-300 NBTR NOOPS D-180
48088 180/1803 00-301 NBTR NOOPS D-181
48093 181/0241 00-302 AXIS-1 Power OFF
48094 181/0422 00-301 NBTR PLAYBACK
48095 181/0520 00-301 NBTR NOOPS D-181
48104 181/1922 00-303 NBTR NOOPS D-182
48109 182/0341 00-303 NBTR PLAYBACK
48110 182/0522 00-303 NBTR NOOPS D-182
48110 182/0522 00-304 HALOE Power OFF (Low S/C Pwr)
48118 182/1805 00-305 NBTR NOOPS D183
________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 30 June 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI HRDI QB CURR MONITOR 48082
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 30 June 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
-None this report period
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 30 June 2000
23 June 2000 - GMT Day 175 Beta = 31.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- ----------- ------ ------- ------
1 +39.2/+11.2 -3.05/-4.31 +0.0/+0.0 15.7 0.000 0.0
2 +16.8/-56.0 +5.47/+4.51 +28.8/-14.4 25.1 1.040 20.7
3 +22.4/-28.0 +4.51/+3.23 +28.0/-14.4 25.1 1.037 20.5
30 June 2000 - GMT Day 182 Beta = deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------ ------- ------
-Data not available at this time
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 30 June 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-028 FOT 48035 177/0555 22574 Premature LOS
00-029 FOT 48037 177/0838 22575 Premature LOS
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-027 FOT 47928 170/0215 22524 Late Acq./Data Loss
00-028 FOT 48035 177/0555 22574 Premature LOS
00-029 FOT 48037 177/0838 22575 Premature LOS
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
00-026 FOT see list June 00 14170 Generic Late Acq.
SEVNTFW00/26